### v2

# GraphPFN: A Prior-Data Fitted Graph Foundation Model

Abstract
Graph foundation models face several fundamental challenges including transferability across datasets and data scarcity, which calls into question the very feasibility of graph foundation models. However, despite similar challenges, the tabular domain has recently witnessed the emergence of the first successful foundation models such as TabPFNv2 and LimiX. Many of these models are based on the prior-data fitted networks (PFN) framework, in which models are pretrained on carefully designed synthetic datasets to make predictions in an in-context learning setting. Recently, G2T-FM has made the first step towards adopting PFNs for graphs, yet it is limited to hand-crafted features and was never pretrained on graph data. In this work, we make the next step by proposing GraphPFN, a PFN-based model designed and pretrained specifically for graph node-level tasks. Following the PFN framework, we first design a prior distribution of synthetic attributed graphs by using a novel combination of multi-level stochastic block models and a preferential attachment process for structure generation and graph-aware structured causal models for attribute generation. Then, we augment the tabular foundation model LimiX with attention-based graph neighborhood aggregation layers and train it on synthetic graphs sampled from our prior. On diverse real-world graph datasets with node-level tasks, GraphPFN shows strong in-context learning performance and achieves state-of-the-art results after finetuning, outperforming both G2T-FM and task-specific GNNs trained from scratch on most datasets. More broadly, GraphPFN shows the potential of PFN-based models for building graph foundation models. Our code is available at https://github.com/yandex-research/graphpfn.
1 Introduction
In recent years, foundation models have significantly advanced the state of the art in domains such as natural language processing and computer vision. These models can learn from large unannotated datasets and generalize across a wide range of downstream tasks. Notable early examples include BERT (Devlin et al., 2019) and GPT-3 (Brown et al., 2020) in natural language processing, as well as ViT (Dosovitskiy et al., 2020) and CLIP (Radford et al., 2021) in computer vision. These models often use self-supervised or unsupervised pretraining to learn rich, transferable representations. This approach has changed how models are built and used, removing the need for task-specific models and reducing the dependence on large labeled datasets for every single task. Inspired by these successes, there is growing interest now in extending the foundation models methodology to other modalities, including graphs.
However, developing graph foundation models (GFMs) is much more challenging. Unlike text and images, graph data does not constitute a single domain. Instead, graphs are used to represent data from different domains, e.g., social networks (both virtual and real-world), information networks, transportation networks, co-purchasing networks, various physical, biological, or engineering systems, or even networks of abstract concepts. As a consequence, both the structure of a graph and its attributes (features and labels) may vary significantly across graph datasets and tasks. Moreover, the amount and diversity of the available graph data are significantly lower compared to those in computer vision and natural language processing. Taken together, these challenges call into question the very feasibility of graph foundation models.
However, despite facing similar challenges, tabular machine learning has recently witnessed the emergence of the first successful tabular foundation models (TFMs) such as TabPFNv2 (Hollmann et al., 2025) or LimiX (Zhang et al., 2025a). Many of these models are based on the framework of prior-data fitted networks (PFNs, Müller et al., 2022; Hollmann et al., 2023). PFNs are pretrained on diverse synthetic datasets drawn from a prior to approximate Bayesian inference and can make predictions via in-context learning. Strong performance in the tabular domain suggests that PFNs offer a promising path towards building foundation models for both tabular and graph domains.
The recent works G2T-FM (Eremeev et al., 2025), TAG (Hayler et al., 2025), and TabPFN-GN (Choi et al., 2025) have made the first step towards adopting PFNs for graph tasks by utilizing foundation models for tabular data to create graph foundation models for node-level tasks. For this, they augment node features with graph-based information such as neighborhood-aggregated features or Laplacian positional encodings. This allows for transforming a graph node-level prediction problem into a tabular prediction problem and applying an existing tabular foundation model to this task. The resulting approach shows strong performance, but such models still depend heavily on hand-crafted features and lack large-scale pretraining on diverse graph data. As a result, they are limited in their ability to capture complex graph patterns.
In this work, we make the next step by proposing GraphPFN, a PFN-based model designed and pretrained specifically for graph node-level tasks. Following the PFN framework, we pretrain GraphPFN on synthetic datasets drawn from a carefully designed graph prior. For generating graph structures, we propose an approach that combines multiple stochastic block models and augments them with a preferential-attachment process. We then generate graph-structure-dependent node attributes for our graphs by augmenting tabular structured causal models (SCMs, Hollmann et al., 2023; Qu et al., 2025) typical for tabular PFNs with message-passing mechanisms at random SCM nodes. This method allows us to efficiently generate millions of realistic and diverse synthetic graph datasets. Then, we initialize GraphPFN from the tabular foundation model LimiX (Zhang et al., 2025a) and add an attention-based message-passing layer to each block of LimiX. This allows the model to learn complex graph-specific patterns while retaining its ability to handle diverse features and labels inherited from LimiX.
Our experiments show that on diverse real-world graph node-level prediction datasets, GraphPFN achieves strong in-context learning performance, competitive with the best current models — well-tuned traditional GNNs (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Shi et al., 2021) with improved architectures (Platonov et al., 2023b) and recent TFM-based GFMs G2T-FM (Eremeev et al., 2025) and TAG (Hayler et al., 2025). Furthermore, after finetuning, GraphPFN outperforms all other approaches, setting a new state of the art for the considered datasets.
Overall, our main contributions can be summarized as follows:
  • We propose GraphPFN, which is, to the best of our knowledge, the first publicly available PFN-based model designed and pretrained specifically for graph node-level tasks.
  • We introduce a novel graph prior for the efficient generation of realistic synthetic attributed graphs.
  • We demonstrate that the finetuned GraphPFN outperforms both strong traditional GNN baselines and existing graph foundation models.
More generally, our study shows that pretraining graph-aware PFNs on synthetic graph datasets from a well-designed graph prior is a promising direction for building powerful and generalizable graph foundation models.
2 Related work
2.1 Prior-data fitted networks
Prior-data fitted networks (PFNs) were first introduced by Müller et al. (2022). The main idea behind PFNs is to train models that can make predictions on previously unseen datasets in a single forward pass. These models leverage in-context learning (ICL): rather than updating model parameters for each new dataset, they use the context provided at inference time to adapt their predictions without additional training.
In the PFN framework, the input to the model consists of two parts: a set of training samples with their labels, called the context, and a set of test samples without labels, called the query. During a single forward pass, the model uses the context to make predictions for the query samples, thus performing ICL. In practice, PFNs are commonly implemented as Transformers with a specific attention mask (Hollmann et al., 2023; 2025): context (training) samples attend to all other context samples, while query (test) samples are only allowed to attend to context samples and not to each other. This structure ensures that predictions for each query are based solely on the training data.
PFNs are trained via pretraining on a large collection of synthetic datasets. To achieve this, one specifies a prior over supervised datasets, and the model is trained to perform ICL as described above, predicting labels for query samples given context samples. As shown by Müller et al. (2022), this procedure trains the network to approximate the posterior predictive distribution under the chosen prior, which provides the main theoretical motivation for the approach.
2.2 Tabular foundation models
In their pioneering work TabPFN (Hollmann et al., 2023) and its successor TabPFNv2 (Hollmann et al., 2025), the authors proposed to utilize the framework of prior-data fitted networks to create tabular foundation models and showed that such models can achieve strong results, competitive with other approaches (Erickson et al., 2025). Nowadays, TFMs have become an active research area, with several new methods released recently (Zhang et al., 2025a; b; Grinsztajn et al., 2025; Qu et al., 2026). Some methods focus on scalability (Qu et al., 2025) or faster inference (Mueller et al., 2025), while others emphasize training on real-world datasets rather than synthetic tasks (Ma et al., 2025). Together, these works broaden the design space of tabular foundation models by trading off data sources, computational efficiency, and scalability.
2.3 Graph foundation models
Similar to foundation models for tabular data, graph foundation models face the challenge of handling datasets from diverse domains. A particularly difficult, yet crucial, aspect is managing the wide variety of node attributes (features and labels) present in different graphs. Early GFMs did not fully address this issue. They often relied on dimensionality reduction techniques such as principal component analysis or singular value decomposition (Xia & Huang, 2024; Zhao et al., 2024; Wang et al., 2025; Yu et al., 2025), or restricted their focus to graphs where node attributes are all textual (Wang et al., 2024; He et al., 2025; Liu et al., 2024).
More recent works, such as G2T-FM (Eremeev et al., 2025), TAG (Hayler et al., 2025), and TabPFN-GN (Choi et al., 2025), have explored leveraging tabular foundation models to better address feature diversity in graph datasets. These approaches incorporate hand-crafted features, for example, neighborhood-aggregated features or Laplacian positional encodings, to effectively convert graph information into tabular features. Empirical results show that these methods achieve strong results, always significantly outperforming prior GFMs (Xia et al., 2024; Xia & Huang, 2024; Zhao et al., 2024; Finkelshtein et al., 2025; Zhao et al., 2025) and frequently outperforming well-tuned GNNs trained from scratch (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Shi et al., 2021) with improved architectures (Platonov et al., 2023b), supporting the utility of employing tabular foundation models as a basis for learning on graph data.
3 GraphPFN
GraphPFN is a foundation model designed for in-context learning on graph-structured data. Inspired by recent advances in prior-data fitted networks (PFNs) for tabular data (Hollmann et al., 2025), GraphPFN extends these ideas to graphs by augmenting a tabular foundation model with attention-based message-passing adapters. This design allows GraphPFN to reuse strong feature modeling from tabular pretraining while capturing complex graph-specific patterns.
3.1 Architecture
Our model architecture extends the tabular foundation model LimiX (Zhang et al., 2025a) by adding attention-based message-passing layers to each of its transformer blocks. This strategy is inspired by recent findings that tabular foundation models can already learn patterns relevant for various graph tasks (Eremeev et al., 2025; Hayler et al., 2025; Choi et al., 2025). By initializing our model with a pretrained tabular foundation model instead of training from scratch, we leverage these learned representations, which significantly reduces computational costs while still achieving strong results.
Below, we first summarize the common architecture of a tabular foundation model (Hollmann et al., 2025; Zhang et al., 2025a), considering LimiX as the specific example, to clarify how GraphPFN represents samples (nodes) and features, and how attention flows in the base model. We then describe how the graph adapters modify this flow to leverage the graph topology.
(a) Different attention mechanisms in GraphPFN.
(b) Base neural block in GraphPFN architecture.
Figure 1: An illustration of GraphPFN architecture.
LimiX
LimiX (Zhang et al., 2025a) is a transformer-style foundation model for tabular data that departs from the common design of representing each sample with a single fixed-length embedding. Instead, it uses a multi-token representation: for every sample, each feature contributes one token, [^1] In the current implementation, two features are grouped into one token, but we omit this detail for clarity. yielding a token grid with one axis for features and one for samples. This design naturally handles a variable number of heterogeneous features in different datasets without changing and retraining the model.
A LimiX transformer block contains three attention layers, each followed by an element-wise feed-forward network (FFN). Two layers are feature-level multi-head attention (MHA) modules that operate within a sample, allowing all feature tokens of the same sample to attend to each other. The third is a sample-level MHA that operates within a feature across samples, allowing tokens corresponding to the same feature to exchange information across the dataset. The feature-level MHAs enable rich interactions among features within each sample, while the sample-level MHA supports in-context learning by transferring information across samples for the same feature.
Attention masking at the sample level follows the standard PFN protocol: training (context) samples attend to all other training samples, and test (query) samples attend only to training samples. Thus, information can flow from train to test but not from test to train or between test samples. Feature-level attention within a sample is unmasked.
Graph adapters
To inject graph structure information without disrupting LimiX’s tokenization, we add a message-passing adapter that implements scaled dot product attention between neighboring nodes at the end of every LimiX transformer block (see Figure 1 for an illustration). Note that we use the scaled dot product attention that is identical to the original Transformer attention (Vaswani et al., 2017) except for being restricted to 1-hop graph neighborhoods. Intuitively, our message-passing adapter performs a second, graph-structure-aware round of sample-level attention: tokens may exchange information only along the observed edges. In contrast to the global sample-level attention of LimiX (which is masked by the PFN protocol), the graph adapter is masked by the adjacency and therefore routes information locally, from each node to its neighbors. Because we keep the per-feature token representation intact, the adapter runs over the sample axis for each feature token independently, using the same graph mask across features.
This design complements the global PFN-style attention by adding a local channel that is common in graph learning. We process the entire graph jointly and the graph adapter allows bidirectional exchange between labeled and unlabeled nodes along edges. Similar to the classic GNNs, these message-passing layers allow the model to capture complex graph dependencies that cannot be captured by hand-crafted features.
Each adapter is implemented as a sparse, multi-head attention module with the adjacency as its mask, where two nodes attend to each other if and only if an edge connects them. Similar to other attention modules, it is followed by a feed-forward network independently applied to each token. Both the FFN and the message-passing layer are wrapped with residual connections (He et al., 2016) and layer normalization (Ba et al., 2016), mirroring the structure of the LimiX blocks for stable optimization.
3.2 Pretraining
GraphPFN was pretrained under the PFN framework (Müller et al., 2022; Hollmann et al., 2023; 2025) by continuing training from the LimiX checkpoint. In total, we used 2,240,0002{,}240{,}000 synthetic datasets generated according to the prior described below. Pretraining ran for approximately 7 days on 8 NVIDIA A100 80GB GPUs. Each GPU processed one synthetic dataset per step with 20 gradient accumulation steps, resulting in 160 datasets per optimizer step.
Inspired by TabICL (Qu et al., 2025), we pretrained GraphPFN in two stages to optimize training efficiency by gradually increasing the size of synthetic datasets. Specifically, Stage 1 ran for 10,00010{,}000 optimizer steps with sample sizes drawn log-uniformly from 1,0001{,}000 to 2,0002{,}000 and took approximately 2 days. Stage 2 ran for 4,0004{,}000 optimizer steps with sample sizes drawn log-uniformly from 8,0008{,}000 to 10,00010{,}000 and took approximately 5 days.
We optimized the model using AdamW (Loshchilov & Hutter, 2019) with weight decay $0.1$. We used a cosine annealing learning rate schedule (Loshchilov & Hutter, 2017) with linear warmup over the first $10\%$ of training steps and a base learning rate of $3\cdot 10^{-4}$. To improve training stability and preserve the feature modeling capabilities of LimiX, we froze all model layers except the graph adapters. Only these components were updated during pretraining. To further stabilize training, we applied an exponential moving average (EMA) to the weights throughout pretraining, with decay $0.98$ in Stage 1 and $0.95$ in Stage 2.
Objective
We optimize a joint objective that combines the PFN supervised loss with the masked graph modeling (MGM) loss (Li et al., 2023). For the supervised PFN term, we sample a random set of context nodes for each dataset, compute predictions for all other nodes, and minimize the supervised loss on them. We use the cross-entropy loss for classification tasks and the mean squared error loss for regression tasks.
To encourage more graph-aware representations, inspired by the success of GNN self-supervised pretraining, we add the masked graph modeling term from Li et al. (2023). Specifically, we randomly sample a fraction $p=0.1$ of edges as positive samples, remove these edges from the input graph, and uniformly sample an equal number of unconnected node pairs as negative examples. We then train the model with the cross-entropy loss to distinguish between positive and negative edges. For this, we apply an additional MLP head to the pointwise multiplication of the target embeddings from the last layer for the source and destination nodes of each sampled edge. The total loss is the sum of the supervised loss and the MGM loss, with a coefficient of $0.1$ applied to the MGM term.
Learning curves
To monitor the progress of pretraining, we evaluated the in-context learning performance of GraphPFN on several datasets from the GraphLand benchmark (Bazhenov et al., 2025) (see Section 5.1 for the evaluation procedure) every 100 steps during the first stage and every 20 steps during the second stage. We emphasize that these evaluations were used only for a post-hoc analysis of training progress: test set performance was not used for early stopping or any other form of model selection. Figure 2 shows that the first stage enables GraphPFN to quickly reach reasonable performance on most datasets. On datasets such as artnet-views, tolokers-2, and twitch-views, where the performance gap between the graph-agnostic LimiX and GraphPFN is largest, most of the gap is closed within the first 1-3K gradient steps. Nevertheless, the second stage consistently improves performance across all datasets and is essential for achieving strong final performance of GraphPFN.
Figure 2: In-context learning performance of intermediate checkpoints of GraphPFN. The $x$-axis represents the number of steps, and the $y$-axis represents the metric on the test set. We also add vertical line at 10K steps to separate two pretraining stages.
4 Graph prior
As discussed above, GraphPFN is based on the prior-data fitted networks (PFNs) framework (Müller et al., 2022). In this approach, the model is pretrained on a large number of synthetic graph datasets sampled from a chosen prior distribution. Since the pretrained model aims to approximate the posterior predictive distribution, it is crucial to design a diverse, high-quality, and realistic prior. In this section, we describe the prior used for pretraining GraphPFN. First, we explain our method for generating realistic graph structures. Then, we describe how we use these graphs to generate node attributes and targets.
4.1 Structure generation
Our main aim is to generate graph structures similar to real-world graphs. After examining graphs from a range of graph machine learning datasets, we find that most of them exhibit strong clustering (community) structure, which in general is a common feature of real-world graphs (Girvan & Newman, 2002). Thus, as the basis for our graph generation process, we use the degree-corrected stochastic block model (SBM) (Karrer & Newman, 2011), which can generate graphs with community structure. However, we find that graphs generated from the degree-corrected SBM exhibit clusters that appear overly regular and well-defined, resembling separated spheres. In contrast, clusters in real-world graphs are often more irregular, with more complex shapes and frequent overlaps. Thus, to obtain graph structures similar to real-world ones, we design a novel method that combines multiple SBMs. First, we generate several first-level graphs from SBMs with different parameters. Then, we generate a second-level graph from another SBM, such that this second-level graph has the number of nodes equal to the total number of nodes across all first-level graphs. We then randomly assign each node from the first-level graphs to a unique node in the second-level graph, thus essentially constructing a bijection $f$ between first-level and second-level graph nodes. Then, we transfer each edge from the first-level graphs to the second-level graph by creating a new edge in the second-level graph between the corresponding nodes, i.e., if there was an edge between nodes $u$ and $v$ in the first-level graphs, then we create an edge between nodes $f(u)$ and $f(v)$ in the second-level graph. The obtained second-level graph with additional edges combines multiple graphs generated from different SBMs and exhibits clusters of nodes with complex shapes and overlaps that we aimed to capture.
Further, we observe that most graphs from graph benchmarks exhibit a core-periphery structure, where the core is composed of multiple relatively dense clusters, but there are also many low-degree peripheral nodes. Such a structure is known to be common in real-world networks (Zhang et al., 2015). While our method of combining multiple graphs generated from SBMs produces realistic node clusters, it produces a relatively small number of peripheral nodes. Thus, we augment the approach discussed above with a preferential attachment (PA) process (Price, 1965; 1976; Albert & Barabási, 2002). Specifically, we use the graph obtained thus far as the initialization for the PA process. We sequentially add low-degree nodes and connect them to the previous ones with probabilities proportional to their degrees. The initial degree of each new node is chosen randomly.
Our method has many hyperparameters such as the number and size of blocks for SBMs or their degree sequences. Similar to prior works on PFNs (Hollmann et al., 2023; 2025; Qu et al., 2025; 2026), we define probability distributions for each hyperparameter and sample new hyperparameters for each synthetic graph, which allows us to generate diverse graphs. At the same time, we can easily set bounds for sizes, densities, or maximum degrees of the generated graphs, allowing us to ensure that the graphs fit the desired constraints.
We provide example visualizations of several graphs generated by our process in Appendix B.1. We also analyze how training on simpler random graphs affects the downstream performance (Appendix B.3). Finally, we inspect the alignment of various characteristics between real and synthetic datasets in Appendix B.2.
4.2 Attribute generation
We generate features and targets for synthetic graphs with a neural structural causal model (SCM) that extends the MLP-based SCM of Qu et al. (2025); Hollmann et al. (2023). As a starting point, we follow the TabICL protocol: we sample an MLP architecture (number of layers, dimension of hidden layers, activation function) and its weights at random, draw random inputs, propagate them through the network, and then designate a random subset of neurons as observed features and another random neuron as the target, leaving the rest as latent variables. This yields a broad family of causal mechanisms in which features and targets can depend on each other and on latent confounders. We refer to Qu et al. (2025); Hollmann et al. (2023) for further details.
To make attributes also depend on the graph structure, we extend this SCM in two complementary ways. First, we introduce a mixture of MLP and GNN neurons. For each dataset, we sample a mixing probability $p\in\{0.0,0.1,\ldots,0.9,1.0\}$ and a GNN type from $\{\mathrm{GCN},\mathrm{SAGE_{mean}},\mathrm{SAGE_{max}},\mathrm{SAGE_{min}},\mathrm{GT}\}$. At every hidden layer, we compute the outputs of both an MLP transformation and a GNN layer. Each neuron is then independently assigned to be MLP-type or GNN-type, with probabilities 1p1-p and $p$, respectively, and its value is taken from the corresponding transformation. This mechanism controls how strongly the generated variables depend on the graph. Second, with a $0.5$ probability, we augment the random inputs with Laplacian positional encodings (LapPE) (Dwivedi et al., 2020; Belkin & Niyogi, 2001) of dimension uniformly sampled from $\{1,\ldots,32\}$ to further integrate graph structure into the data generation process.
Together, the mixed MLP/GNN neurons and optional LapPE inject graph information into the SCM while remaining close to the tabular prior. When $p=0$ and LapPE is not used, the procedure reduces to the TabICL-style tabular SCM, while a larger value of $p$ and the inclusion of LapPE increase the influence of graph structure on both features and targets.
5 Experiments
5.1 Experimental setup
Datasets
In terms of dataset selection, we follow the experimental setup of G2T-FM (Eremeev et al., 2025). We evaluate two collections of datasets: (i) real-world datasets from the recently proposed GraphLand benchmark (Bazhenov et al., 2025); and (ii) some of the classic graph datasets. Together, these datasets cover node classification and regression, come from diverse application domains, include both homophilous and non-homophilous graphs, [^2] A graph is called homophilous if its edges tend to connect nodes with similar labels, see Newman (2003); Platonov et al. (2023a); Mironov & Prokhorenkova (2024) for details. and span a range of densities and other graph structural properties. Table 7 lists the datasets used in our study and summarizes their statistics. For all datasets, we use 10%/10%/80% train/validation/test splits. Due to current limitations of TFMs, we restrict our study to small- and medium-scale datasets and exclude classification tasks with more than 10 classes. [^3] In principle, TFMs can handle more than 10 classes via schemes such as error-correcting output codes, as proposed in Hollmann et al. (2025). For simplicity, we focus on datasets with at most 10 classes that are natively supported by existing TFMs.
In our evaluation, we run all experiments 10 times and report the mean and standard deviation of the model performance. We report average precision for binary classification tasks, accuracy for multiclass classification tasks, and $R^{2}$ for regression tasks. For all metrics, higher is better.
Methods
In addition to the proposed GraphPFN, we evaluate the following methods:
  • LightGBM (Ke et al., 2017), a strong tabular baseline, augmented with neighborhood feature aggregation (NFA) (Bazhenov et al., 2025) to incorporate information about the graph structure.
  • Classic GNNs: GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), neighborhood-attention Graph Transformer (GT) (Shi et al., 2021). Following Platonov et al. (2023b), we augment these models with residual connections (He et al., 2016), layer normalization (Ba et al., 2016), and MLP blocks, which have been shown to substantially improve the performance of classic GNNs (Luo et al., 2024; 2025). We perform extensive hyperparameter tuning for these models.
  • Prior GFMs: OpenGraph (Xia et al., 2024), AnyGraph (Xia & Huang, 2024), GCOPE (Zhao et al., 2024), and TS-GNN (Finkelshtein et al., 2025), which are not based on the PFN framework.
  • Recent PFN-based GFMs: G2T-FM (Eremeev et al., 2025) and TAG (Hayler et al., 2025). To the best of our knowledge, these are the strongest publicly available graph foundation models for node-level tasks with arbitrary features.
We evaluate GraphPFN in both in-context learning (ICL) and full finetuning (FT) settings. Other models are evaluated in ICL, FT, or both settings based on their official implementations. For G2T-FM and GraphPFN in the ICL setting, we additionally report ensembling (E) results obtained by averaging the predictions over 10 forward passes with different random seeds. Since TAG inherently employs ensembling, we also mark it with “(ICL, E)”, despite the exact ensembling implementation being different from G2T-FM and GraphPFN. For the FT setting, we finetune the entire model, following Eremeev et al. (2025); Rubachev et al. (2025). Inspired by Abboud et al. (2020); Sato et al. (2021); Kanatsoulis et al. (2025), we additionally augment the input of GraphPFN with 8 random features during each forward pass in all experiments in this section. For further details on the experimental setup, please refer to Appendix D.
Table 1: Evaluation results on the GraphLand datasets under the RL (Random Low) data split. ICL stands for in-context learning, FT stands for fine-tuning, and E stands for ensembling. We report average precision for binary classification tasks and $R^{2}$ for regression tasks. We also report average rank over classification datasets in the “AR (cls)” column and over all datasets in the “AR (all)” column. During the computation of “AR (all)” we exclude all methods that do not support regression tasks, so that ranks are computed only over tables with no missing entries. For each column, we highlight first, second, and third best results with a color.
artnet-exp city-reviews tolokers-2 artnet-views avazu-ctr city-roads-M hm-prices twitch-views AR (cls) AR (all)
LightGBM + NFA $46.13\pm 0.04$ $78.53\pm 0.01$ $56.34\pm 0.06$ $56.10\pm 0.02$ $31.71\pm 0.01$ $61.18\pm 0.03$ $70.84\pm 0.04$ $60.14\pm 0.01$ $\mathbf{11.0}$ $\mathbf{9.75}$
GCN $44.86\pm 0.36$ $77.81\pm 0.15$ $56.27\pm 0.31$ $56.03\pm 0.25$ $32.00\pm 0.16$ $58.82\pm 0.25$ $68.02\pm 0.42$ $75.51\pm 0.05$ $\mathbf{13.33}$ $\mathbf{10.25}$
GraphSAGE $45.14\pm 0.36$ $78.17\pm 0.10$ $54.43\pm 0.34$ $49.32\pm 0.91$ $31.44\pm 0.16$ $59.44\pm 0.27$ $70.00\pm 0.74$ $66.29\pm 0.32$ $\mathbf{12.33}$ $\mathbf{11.25}$
GAT $45.06\pm 0.52$ $77.74\pm 0.21$ $57.41\pm 0.85$ $53.60\pm 0.24$ $32.63\pm 0.17$ $59.86\pm 0.20$ $72.07\pm 1.22$ $72.89\pm 0.27$ $\mathbf{12.67}$ $\mathbf{9.5}$
GT $46.41\pm 0.71$ $77.34\pm 0.21$ $56.98\pm 0.55$ $53.37\pm 0.46$ $31.11\pm 0.49$ $59.55\pm 0.28$ $69.44\pm 0.94$ $72.13\pm 0.13$ $\mathbf{12.67}$ $\mathbf{10.88}$
OpenGraph (ICL) $15.16\pm 0.83$ $59.09\pm 0.72$ $40.38\pm 1.13$ $\mathbf{17.0}$
AnyGraph (ICL) $12.84\pm 0.93$ $63.71\pm 1.45$ $28.75\pm 3.56$ $\mathbf{18.33}$
TS-GNN (ICL) $20.44\pm 1.05$ $43.46\pm 5.17$ $38.54\pm 0.94$ $\mathbf{17.33}$
G2T-LimiX (ICL) $48.44\pm 0.23$ $77.81\pm 0.49$ $61.45\pm 0.33$ $60.94\pm 0.11$ $32.39\pm 0.13$ $64.53\pm 0.09$ $74.96\pm 0.07$ $71.08\pm 0.07$ $\mathbf{7.33}$ $\mathbf{7.5}$
GraphPFN (ICL) $50.85\pm 0.31$ $79.89\pm 0.10$ $60.33\pm 0.60$ $61.77\pm 0.23$ $30.41\pm 0.18$ $63.55\pm 0.29$ $75.38\pm 0.26$ $73.27\pm 0.13$ $\mathbf{5.67}$ $\mathbf{7.12}$
G2T-LimiX (ICL, E) $48.52\pm 0.29$ $77.97\pm 0.53$ $61.58\pm 0.32$ $60.96\pm 0.10$ $32.41\pm 0.14$ $64.69\pm 0.05$ $74.97\pm 0.06$ $71.30\pm 0.07$ $\mathbf{6.33}$ $\mathbf{6.5}$
TAG-TabPFNv2 (ICL, E) $47.87\pm 0.39$ $77.38\pm 0.29$ $59.33\pm 1.00$ $\mathbf{11.33}$
TAG-LimiX (ICL, E) $50.19\pm 0.46$ $78.87\pm 0.13$ $59.46\pm 1.05$ $\mathbf{7.0}$
GraphPFN (ICL, E) $51.49\pm 0.10$ $80.22\pm 0.04$ $60.98\pm 0.25$ $62.07\pm 0.04$ $30.46\pm 0.14$ $64.29\pm 0.12$ $75.82\pm 0.06$ $73.81\pm 0.04$ $\mathbf{3.0}$ $\mathbf{5.25}$
GCOPE (FT) $14.92\pm 1.56$ $67.16\pm 0.98$ $28.81\pm 1.28$ $\mathbf{17.33}$
G2T-LimiX (FT) $49.89\pm 0.18$ $80.14\pm 0.06$ $61.28\pm 0.70$ $62.08\pm 0.12$ $34.02\pm 0.27$ $65.82\pm 0.18$ $76.27\pm 0.24$ $73.69\pm 0.35$ $\mathbf{5.0}$ $\mathbf{4.25}$
GraphPFN (FT) $50.64\pm 1.27$ $80.49\pm 0.16$ $60.90\pm 0.98$ $63.90\pm 0.12$ $34.34\pm 0.50$ $66.04\pm 0.41$ $79.26\pm 0.43$ $78.17\pm 0.13$ $\mathbf{4.0}$ $\mathbf{2.75}$
G2T-LimiX (FT, E) $49.96\pm 0.09$ $80.15\pm 0.07$ $60.71\pm 0.57$ $62.11\pm 0.06$ $33.86\pm 0.33$ $65.96\pm 0.11$ $76.30\pm 0.19$ $73.62\pm 0.49$ $\mathbf{5.67}$ $\mathbf{4.38}$
GraphPFN (FT, E) $51.97\pm 0.40$ $80.89\pm 0.05$ $60.92\pm 0.56$ $65.12\pm 0.11$ $35.05\pm 0.23$ $67.27\pm 0.14$ $80.13\pm 0.37$ $78.50\pm 0.36$ $\mathbf{2.33}$ $\mathbf{1.5}$
5.2 Experimental results
Tables 1 and 2 present the results of our experiments. Below, we summarize and discuss our key observations. Additional experimental results are provided in Appendix C.
Observation 1
GraphPFN shows strong ICL performance, outperforming GNNs and other GFMs on the GraphLand benchmark, while matching them on classic graph datasets.
Specifically, when compared with GNNs and other GFMs evaluated in an ICL regime, GraphPFN (ICL) achieves the best average rank on GraphLand datasets and only loses to GAT on classic datasets. Importantly, unlike G2T-FM, which also achieves strong ICL performance, GraphPFN does not require potentially heavy preprocessing like the computation of Laplacian positional encodings or PageRank.
Observation 2
Ensembling further boosts the ICL performance of GraphPFN, in particular allowing it to outperform GNNs on average on both collections of datasets.
Specifically, GraphPFN (ICL, E) achieves better average rank than GNNs and G2T-LimiX (ICL, E) on both collections of datasets. When compared to TAG, GraphPFN outperforms it on all considered GraphLand datasets, and performs on par on classic datasets, while requiring no potentially heavy preprocessing.
Observation 3
Finetuning allows GraphPFN to achieve the best results on 10 out of 13 datasets, and the gains over the second-best method are often substantial.
Specifically, the finetuned and ensembled GraphPFN achieves the best performance among all considered methods on all datasets except tolokers-2, amazon-ratings, and pubmed. Moreover, GraphPFN often yields substantial gains over the second-best method, for example, of at least one percentage point on the artnet-exp, artnet-views, avazu-ctr, city-roads-M, hm-prices, twitch-views, and questions datasets.
The performance of finetuned GraphPFN without ensembling is also strong. Specifically, it achieves the second-best result across all considered methods on 7 out of 13 datasets, only losing to GraphPFN (FT, E). It also achieves better average ranks than all other methods except those evaluated in the (FT, E) regime on both collections of datasets.
We hypothesize that this strong performance stems from GraphPFN’s ability to capture complex graph patterns via message passing, in contrast to G2T-FM and TAG, which rely on hand-crafted graph-based features. At the same time, the pretraining procedure appears to be critical. When we replace the pretrained graph adapters with randomly initialized ones and finetune on a downstream dataset, performance drops substantially; see Appendix C for details.
Table 2: Evaluation results on classic graph datasets under a random $10\%/10\%/80\%$ train/val/test split. ICL stands for in-context learning, FT stands for fine-tuning, and E stands for ensembling. We report average precision for binary classification tasks and accuracy for multiclass classification tasks. We also report average rank (AR) in the last column. For each column, we highlight first, second, and third best results with a color.
amazon-ratings facebook pubmed questions wiki-cs AR
GCN $41.43\pm 0.49$ $91.26\pm 0.21$ $85.46\pm 0.19$ $15.42\pm 0.67$ $81.74\pm 0.21$ $\mathbf{11.4}$
GraphSAGE $40.07\pm 0.53$ $91.12\pm 0.22$ $86.04\pm 0.27$ $16.55\pm 0.64$ $81.50\pm 0.27$ $\mathbf{11.6}$
GAT $40.67\pm 0.55$ $92.61\pm 0.21$ $84.81\pm 0.24$ $16.75\pm 0.67$ $82.25\pm 0.27$ $\mathbf{9.2}$
GT $41.56\pm 0.40$ $91.71\pm 0.22$ $84.95\pm 0.19$ $14.03\pm 0.90$ $82.54\pm 0.21$ $\mathbf{10.4}$
OpenGraph (ICL) $29.36\pm 1.24$ $75.27\pm 5.05$ $70.30\pm 2.67$ $3.77\pm 0.65$ $75.66\pm 0.39$ $\mathbf{16.8}$
AnyGraph (ICL) $33.49\pm 3.44$ $61.17\pm 8.64$ $65.31\pm 6.26$ $4.27\pm 0.66$ $65.17\pm 2.51$ $\mathbf{17.0}$
TS-GNN (ICL) $43.00\pm 0.13$ $77.87\pm 2.73$ $64.41\pm 5.11$ $5.00\pm 0.48$ $46.25\pm 9.77$ $\mathbf{15.6}$
G2T-LimiX (ICL) $44.20\pm 0.26$ $91.30\pm 0.23$ $88.93\pm 0.46$ $15.32\pm 0.81$ $80.12\pm 0.23$ $\mathbf{10.2}$
GraphPFN (ICL) $42.09\pm 2.06$ $90.44\pm 0.37$ $89.79\pm 0.35$ $18.27\pm 2.24$ $79.58\pm 0.40$ $\mathbf{10.0}$
G2T-LimiX (ICL, E) $44.65\pm 0.18$ $91.66\pm 0.05$ $89.19\pm 0.18$ $15.46\pm 0.93$ $1.09\pm 0.09 8.8\mathbf${8.8}
TAG-TabPFNv2 (ICL, E) $43.97\pm 0.52$ $93.11\pm 0.17$ $87.80\pm 0.20$ $15.07\pm 1.32$ $82.59\pm 0.12$ $\mathbf{7.6}$
TAG-LimiX (ICL, E) $44.89\pm 0.15$ $92.85\pm 0.14$ $88.09\pm 0.18$ $15.53\pm 1.99$ $82.64\pm 0.17$ $\mathbf{5.8}$
GraphPFN (ICL, E) $44.52\pm 0.27$ $91.70\pm 0.21$ $90.36\pm 0.13$ $20.79\pm 0.49$ $81.69\pm 0.15$ $\mathbf{6.4}$
GCOPE (FT) $39.90\pm 0.43$ $85.08\pm 0.17$ $79.35\pm 0.70$ $6.59\pm 0.43$ $59.13\pm 1.20$ $\mathbf{15.6}$
G2T-LimiX (FT) $44.00\pm 0.92$ $92.17\pm 0.22$ $90.57\pm 0.13$ $20.16\pm 0.56$ $81.90\pm 0.43$ $\mathbf{6.0}$
GraphPFN (FT) $45.34\pm 0.35$ $92.91\pm 0.22$ $89.79\pm 0.35$ $21.85\pm 1.04$ $81.91\pm 0.44$ $\mathbf{4.0}$
G2T-LimiX (FT, E) $46.05\pm 0.23$ $92.53\pm 0.13$ $90.88\pm 0.14$ $20.83\pm 0.44$ $83.11\pm 0.39$ $\mathbf{2.6}$
GraphPFN (FT, E) $45.89\pm 0.53$ $93.94\pm 0.12$ $90.36\pm 0.13$ $22.76\pm 0.55$ $83.53\pm 0.25$ $\mathbf{1.6}$
6 Conclusion
In this work, we propose GraphPFN: a prior-data fitted graph foundation model for node-level tasks. Following the PFN framework, GraphPFN was pretrained on synthetic datasets drawn from a carefully designed prior over attributed graphs to perform predictions in the in-context learning setting. Our experiments show strong results for GraphPFN. Even in the ICL setting, GraphPFN often outperforms classic GNNs and prior ICL GFMs. After finetuning, GraphPFN consistently achieves state-of-the-art results, often bringing substantial improvements over the second-best method. We believe our work shows that pretraining graph foundation models on synthetic datasets drawn from a well-designed prior is a promising direction for developing truly generalizable graph foundation models. Despite promising results, the current implementation of GraphPFN has several limitations, which we discuss in Appendix A.
Appendix A Limitations and future work
Despite promising results, the current implementation of GraphPFN has several limitations that may stimulate future research:
  • GraphPFN is difficult to scale to very large datasets since its current implementation requires processing the entire dataset at once, which leads to significant memory consumption. Developing more scalable graph foundation models can be a promising direction for future work. For example, one can utilize more memory-efficient TFMs (like TabICL (Qu et al., 2025; 2026)) as backbones or combine GraphPFN with sampling methods (Hamilton et al., 2017; Zeng et al., 2020; Zou et al., 2019; Chiang et al., 2019; Zeng et al., 2021; Balin & Çatalyürek, 2023).
  • The proposed graph prior is focused on social and information networks and does not cover graphs from specific domains such as traffic networks. Extending the prior with more diverse random graph models (e.g., geometric graphs (Barthélemy, 2011)) can further improve the results of GraphPFN and make its performance more robust.
  • GraphPFN inherits some limitations from its tabular backbone LimiX. For example, GraphPFN does not natively support more than 10 classes in multiclass classification. This limitation can be alleviated with further development of TFMs or by using approaches such as error-correcting output codes, as proposed in Hollmann et al. (2025).
  • Currently, GraphPFN is limited to node-level prediction tasks and cannot handle link prediction or graph-level tasks (e.g., graph classification or regression). While we have taken a first step towards link-level tasks by adding a masked graph modeling head to GraphPFN during pretraining, the model still heavily relies on the presence of node-level labels to make predictions, so we cannot directly apply GraphPFN to link prediction tasks.
Appendix B Graph prior analysis
We analyze the proposed graph prior from three complementary perspectives. First, we visually inspect the generated graph structures in Appendix B.1. Second, we provide a numerical comparison of graph characteristics to demonstrate the alignment between our synthetic prior and real-world datasets in Appendix B.2. Finally, we examine the effect of the specific graph generation procedure on downstream performance in Appendix B.3.
B.1 Synthetic graph examples
We design our prior to generate graphs that are both realistic and diverse. In Figure 3, we provide examples of our synthetic graphs. Note that these graphs tend to exhibit both community structure and core-periphery structure. Our graph generation process allows us to easily control various graph properties such as graph size or density.
B.2 Numerical analysis of graph characteristics
To further assess how closely our synthetic datasets match the properties of real-world graphs, we perform a numerical analysis of several graph characteristics. Specifically, we sample 1,0001{,}000 synthetic datasets from the prior used in the first pretraining stage and compute the following statistics.
First, we report basic structural measures, including the average degree, the clustering coefficient, and an estimate of the average pairwise distance. Second, to quantify the tendency of adjacent nodes to have similar targets, we compute unbiased homophily (Mironov & Prokhorenkova, 2024) for classification datasets and assortativity (Newman, 2003) for regression datasets. Finally, to evaluate how well the graphs follow the power-law degree distribution, we compute the cumulative degree distribution: for each degree value $k$, we estimate $P_{k}=\mathbb{P}(\deg(u)\geq k)$. We then fit a linear regression to the pairs $(\log k,\log P_{k})$ and report the corresponding $R^{2}$.
Table 4 summarizes the distributions of these characteristics for synthetic datasets, while Table 3 reports the same statistics for real datasets. In most cases, the distribution of characteristics in the synthetic datasets covers that of the real-world datasets.
Table 3: Characteristics of real-world graphs from the GraphLand benchmark.
artnet-exp artnet-views avazu-ctr city-reviews city-roads-M hm-prices tolokers-2 twitch-views
Avg. degree $11.12$ $11.12$ $288.04$ $15.66$ $3.75$ $460.92$ $88.28$ $80.87$
Clustering coefficient $0.03$ $0.03$ $0.24$ $0.26$ $0.00$ $0.27$ $0.23$ $0.02$
Avg. pairwise distance $4.34$ $4.43$ $3.72$ $4.78$ $129.89$ $2.39$ $2.84$ $2.84$
Homophily $0.28$ $0.69$ $0.10$
Assortativity $0.19$ $0.18$ $0.74$ $0.12$ -0.41
Degree power-law $R^{2}$ $0.70$ $0.70$ $0.90$ $0.92$ $0.52$ $0.84$ $0.83$ $0.97$
Table 4: Summary of distribution of characteristics of synthetic graphs for the first pretraining stage.
Mean Std Min Q25 Median Q75 Max
Avg. degree $125.64$ $130.27$ $2.39$ $10.24$ $87.27$ $206.33$ $479.03$
Clustering coefficient $0.3$ $0.23$ $0.0$ $0.08$ $0.28$ $0.49$ $0.88$
Avg. pairwise distance $2.92$ $1.35$ $1.57$ $2.01$ $2.28$ $3.39$ $9.76$
Homophily $0.06$ $0.16$ 0.48-0.48 0.01-0.01 $0.0$ $0.08$ $0.96$
Assortativity $0.05$ $0.13$ 0.21-0.21 0.0-0.0 $0.0$ $0.07$ $0.7$
Degree power-law $R^{2}$ $0.68$ $0.21$ $0.11$ $0.51$ $0.7$ $0.86$ $0.99$
B.3 Pretraining on simpler random graphs
To assess the utility of the proposed graph generation procedure, we conduct an ablation study in which we replace it during pretraining with simpler random graph models. Specifically, we consider three alternatives: the Erdős–Rényi model (Erdős & Rényi, 1959), the degree-corrected SBM (Karrer & Newman, 2011), and the preferential attachment model (Albert & Barabási, 2002) with varying initial degrees. Due to the high computational cost, we run only the first stage of pretraining for all three alternatives. For a fair comparison, we evaluate them against an intermediate GraphPFN checkpoint taken after completing the first pretraining stage.
The results, presented in Table 5, indicate that our proposed procedure yields superior performance in most cases. The only exceptions are the Erdős–Rényi model on the city-roads-M dataset and the degree-corrected SBM on artnet-views. For artnet-views, the difference is small and can plausibly be attributed to stochasticity in the pretraining process. For city-roads-M, we hypothesize that the outcome reflects a mismatch between the structural biases of our procedure, which primarily produces graphs resembling information, social, and web networks, and road networks, which typically lack a pronounced community or core-periphery structure. In this setting, the Erdős–Rényi model, which does not enforce these properties, may be more suitable. Conversely, on most of the datasets, the results for the Erdős–Rényi model are worse than for the more complex approaches, which indicates that a complex and diverse prior is important for GraphPFN pretraining.
Table 5: Results of pretraining GraphPFN on simpler graph models.
artnet-exp artnet-views avazu-ctr city-reviews city-roads-M hm-prices tolokers-2 twitch-views
Erdős–Rényi $46.92$ $56.17$ $25.09$ $78.60$ $63.75$ $66.52$ $45.26$ $55.92$
Degree-corrected SBM $47.27$ $60.87$ $25.68$ $78.71$ $59.60$ $69.86$ $58.29$ $62.69$
Preferential attachment $48.16$ $58.56$ $28.09$ $76.88$ $59.14$ $69.16$ $55.53$ $61.78$
Ours $50.61$ $60.05$ $28.31$ $79.46$ $61.65$ $73.37$ $59.04$ $64.23$
Appendix C Additional results
Table 6: Additional comparison of finetuned GraphPFN with finetuned LimiX with randomly initialized graph adapters (GA).
artnet-exp city-reviews tolokers-2 artnet-views avazu-ctr city-roads-M hm-prices twitch-views
GraphPFN (FT) $50.64\pm 1.27$ $80.49\pm 0.16$ $60.90\pm 0.98$ $63.90\pm 0.12$ $34.34\pm 0.50$ $66.04\pm 0.41$ $79.26\pm 0.43$ $78.17\pm 0.13$
Limix + GA (FT) $46.19\pm 0.21$ $78.94\pm 0.35$ $48.55\pm 1.77$ $53.84\pm 0.59$ $26.78\pm 0.16$ $63.41\pm 0.59$ $69.44\pm 0.48$ $74.01\pm 1.41$
Since GraphPFN achieves strong performance in the finetuning setting, one may hypothesize that the performance comes solely from the powerful graph adapters, but not from the pretraining procedure. To test this hypothesis, we consider the following model. We start with LimiX and add graph adapters, so the architecture exactly matches that of GraphPFN. But instead of using pretrained weights for graph adapters, we employ random weights. In order to stabilize training, we initialized the last layers in all graph adapters with zeros, ensuring that the random initialization does not break the model. After that, we finetune this model following exactly the same protocol as GraphPFN. The results of this model and a comparison with GraphPFN are presented in Table 6. One can see that using random adapters instead of pretrained ones leads to significant drops in performance, supporting the importance of pretraining for achieving the strong performance of the finetuned GraphPFN.
Appendix D Additional details on experimental setup
Data
Table 7 presents key statistics of datasets used in our main experiments. Importantly, the selected datasets cover diverse domains, spanning both homophilic and heterophilic structures, diverse feature types (both tabular and text-based) and tasks (both classification and regression).
Table 7: The key statistics of the considered graph datasets.
name # nodes # edges # features mean degree task # classes homophily feature type
artnet-exp 50,40550{,}405 280,348280{,}348 $75$ $11.1$ cls. 2 no tabular
artnet-views 50,40550{,}405 280,348280{,}348 $50$ $11.1$ reg. no tabular
avazu-ctr 76,26976{,}269 10,984,07710{,}984{,}077 $260$ $288.0$ reg. no tabular
city-reviews 148,801148{,}801 1,165,4151{,}165{,}415 $37$ $15.7$ cls. 2 yes tabular
city-roads-M 57,07357{,}073 107,104107{,}104 $26$ $3.8$ reg. yes tabular
hm-prices 46,56346{,}563 10,730,99510{,}730{,}995 $41$ $460.9$ reg. no tabular
tolokers-2 11,75811{,}758 519,000519{,}000 $16$ $88.3$ cls. 2 no tabular
twitch-views 168,114168{,}114 6,797,5576{,}797{,}557 44 $80.9$ reg. no tabular
amazon-ratings 24,49224{,}492 93,05093{,}050 $300$ $7.6$ cls. 55 no text-based
facebook 22,47022{,}470 170,823170{,}823 $128$ $15.2$ cls. 44 yes text-based
pubmed 19,71719{,}717 44,32444{,}324 $500$ $4.5$ cls. 33 yes text-based
questions 48,92148{,}921 153,540153{,}540 $301$ $6.3$ cls. 2 no text-based
wiki-cs 11,70111{,}701 215,603215{,}603 $300$ $36.9$ cls. 10 yes text-based
PCA
Since some datasets (specifically, amazon-ratings and questions) with text-embedding features have relatively high feature dimensionality, which prevented us from directly finetuning on these datasets, we applied PCA to reduce the feature dimensionality to 64. We did not apply PCA to the pubmed dataset, since, in our preliminary experiments, doing so led to degraded performance. As a result, GraphPFN ran out of memory during finetuning on pubmed, and we therefore report its in-context learning (ICL) performance for this dataset.
Ensembling
For G2T-FM and GraphPFN, we report test-time ensembling results. Specifically, we evaluate the same model under $10$ different random seeds and average the resulting predictions. The diversity among ensemble members is induced by several sources of stochasticity: (1) internal randomness of the TFM model (e.g., random positional encodings), (2) random shuffling of labels and input features, (3) PEARL in G2T-FM, and (4) random features in GraphPFN. In G2T-FM, we do not apply feature shuffling, because it degraded performance on some datasets in our preliminary experiments. When ensembling is combined with finetuning, we apply ensembling at every evaluation step, including validation evaluation for early stopping and the final test evaluation. However, during optimization we use a single randomly sampled ensemble member to compute predictions for each optimizer step.
Figure 3: Example visualizations of graphs from our prior.